# Dynamic Target Attack

### v2

@@ Abstract
Existing gradient-based jailbreak attacks typically optimize an adversarial suffix to induce a fixed affirmative response, e.g.,Sure, here is…”. However, this fixed target usually resides in an extremely low-density region of a safety-aligned LLM’s output distribution conditioned on diverse harmful inputs. Due to the substantial discrepancy between the target and the original output, existing attacks require numerous iterations to optimize the adversarial prompt, which might still fail to induce the low-probability target response from the target LLM. In this paper, we propose Dynamic Target Attack (DTA), a new jailbreaking framework relying on the target LLM’s own responses as targets to optimize the adversarial prompts. In each optimization round, DTA iteratively samples multiple candidate responses directly from the output distribution conditioned on the current prompt, and selects the most harmful response as a temporary target for prompt optimization. In contrast to existing attacks, DTA significantly reduces the discrepancy between the target and the output distribution, substantially easing the optimization process to search for an effective adversarial prompt.
Extensive experiments demonstrate the superior effectiveness and efficiency of DTA: under the white-box setting, DTA only needs $200$ optimization iterations to achieve an average attack success rate (ASR) of over $87\%$ on recent safety-aligned LLMs, exceeding the state-of-the-art baselines by over $15\%$. The time cost of DTA is $\thicksim2$6 times less than existing baselines. Under the black-box setting, DTA uses Llama-3-8B-Instruct as a surrogate model for target sampling and achieves an ASR of $85\%$ against the black-box target model Llama-3-70B-Instruct, exceeding its counterparts by over $25\%$. All code and other materials are available here.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide spectrum of natural language tasks [1, 2, 3, 4]. To steer these capabilities toward beneficial and safe outputs, alignment techniques such as Reinforcement Learning from Human Feedback (RLHF) are employed to instill models with human values [5], which will significantly lower the probability of generating affirmative responses on harmful queries.
@@
Despite the effectiveness of these alignment techniques, most existing gradient-based jailbreak methods [6, 7, 8] could jailbreak safety-aligned LLMs with a common paradigm, i.e., optimizing an adversarial suffix $S$ appended to a harmful prompt $P$ to maximize the probability of the target LLM generating a fixed affirmative response $T_{\text{fixed}}$. However, due to the alignment techniques, $T_{\text{fixed}}$ typically lies in an extremely low-density region of the LLM’s output distribution conditioned on $P+S$ (e.g., $\log p(T_{\text{fixed}}\mid P+S)=-52.63$ in Figure 1), creating a substantial discrepancy between the optimization objective (i.e., $T_{\text{fixed}}$) and the target LLM’s original outputs. As a result, these methods usually require thousands of iterations for optimizing $S^{*}$, while the probability of producing $T_{\text{fixed}}$ conditioned on $P+S^{*}$ remains low (e.g., $\log p(T_{\text{fixed}}\mid P+S^{*})=-39.07$ in Figure 1).
Figure 1: DTA directly samples $T_{\text{sampled}}$ from the LLM, which is more probable than $T_{\text{fixed}}$.
Our approach. We propose Dynamic Target Attack (DTA), a new jailbreaking framework that reframes target selection as an adaptive search over relatively high-density regions of the target LLM’s output space conditioned on the prompt, significantly reducing the discrepancy between the targets and LLM outputs. As shown in Figure 2, DTA (i) directly samples candidate harmful responses from high-density regions of the target LLM’s conditional distribution, (ii) selects the most harmful candidate as a temporary target ($T_{\text{sampled}}$) and performs a few optimization steps, and (iii) re-samples the target from the more harmful conditional distribution after optimization, iterating this sampling–optimization cycle. Due to the reduced discrepancy, the proposed sampling–optimization cycle substantially speeds the optimization process to identify an effective adversarial suffix. For example, in Figure 1, before optimization, the log-probability of generating the sampled response $T_{\text{sampled}}$ conditioned on $P+S$ ($\log p(T_{\text{sampled}}\mid P+S^{*})=-36.68$) is 1.4x higher than that of the fixed target $T_{\text{fixed}}$ ($\log p(T_{\text{fixed}}\mid P+S^{*})$=-52.63). After a few optimization steps, the conditional log-probability improves to 28.03-28.03. To further accelerate convergence, DTA then re-samples from the target LLM conditioned on the harmful prompt and the current updated suffix, which increases the chance of selecting an even more harmful, higher-density target for the next round and thereby further reduces the discrepancy between the target and the model’s output distribution.
@@
We demonstrate that DTA effectively jailbreaks multiple recent safety-aligned LLMs, whether under the white-box setting or the black-box setting. Specifically, under the white-box setting, DTA only needs 200 optimization steps to achieve an average Attack Success Rate (ASR) of over 87% on five recent safety-aligned LLMs, which is over 15% higher than its counterparts. Furthermore, under the black-box setting, DTA effectively achieves an ASR of 85% against Llama-3-70B-Instruct, using Llama-3-8B-Instruct as the surrogate model, achieving an ASR of 85%. Our primary contributions can be summarized as follows:
@@
  • We introduce Dynamic Target Attack (DTA), a novel jailbreak framework that effectively jailbreaks LLMs under both white- and black-box settings. In contrast to existing methods, DTA iteratively samples multiple candidate responses directly from the target model’s output distribution, and selects the most harmful as a temporary target for the current optimization round.
  • We conduct extensive experiments to demonstrate the effectiveness and efficiency of DTA. For example, under the white-box setting, DTA requires only 200 optimization steps to reach an average ASR above $87\%$ across five recent safety-aligned models—exceeding existing baselines by over 15%. Under the black-box setting, DTA still outperforms its counterparts by over 25%.
@@ 2 Background
2.1 Paradigms in White-box Jailbreaking
@@
Existing methods, such as GCG [6] and COLD-Attack [7] are prominent implementations of this paradigm, employing token-level and global-level optimization processes, respectively, to find an effective $S^{*}$.
However, recent researchers have recognized that constraining the optimization to a single, fixed target prefix is overly restrictive and can lead to inefficient optimization. To mitigate this limitation, methods like AdvPrefix [8] are proposed. The core idea of AdvPrefix is to relax the objective by introducing a set of multiple, curated target prefixes, $\mathcal{T}=\{T_{1},T_{2},\cdots,T_{m}\}$. The optimization goal then becomes maximizing the probability of generating any of these targets, often approximated by minimizing the loss of the most likely target in the set:
@@
This multi-target approach provides more flexibility but still relies on a predefined set of external targets. Furthermore, researchers have proposed jailbreaking methods based on Reinforcement Learning (e.g., RLbreaker [9]) and adaptive jailbreaking methods (e.g., llm-adaptive-attacks [10]. However, all existing methods rely on externally-defined objectives, which may not adaptively align with different target LLMs’ natural harmful responses.
@@ 2.2 Language Model Decoding and Sampling Strategies
In practice, most LLMs adopt greedy sampling or random sampling with a low temperature, which forms a narrow output space. A safety-aligned LLM usually generates refusal responses from this narrow space to harmful prompts. However, recent work [11] indicates that a broader output space randomly sampled from a safety-aligned LLM usually contains harmful responses. Inspired by this observation, we propose DTA to exploit the potentially harmful responses randomly sampled from the LLM as targets to optimize the adversarial suffix, aiming to induce similar responses under the modes of greedy sampling or random sampling with a small temperature.
@@ 3 Methodology
3.1 Overview of Dynamic Target Attack
The design of DTA begins from a simple intuition: optimizing the adversarial suffix $S$ toward harmful responses that have relatively high probability in the target model’s conditional output distribution is more efficient than using a fixed, relatively low probability response as the target. Given a harmful prompt $P$, DTA learns an adversarial suffix $S$ such that the target LLM produces a complete harmful response under standard decoding. As shown in Figure 2, DTA proceeds in short exploration–optimization cycles:
  1. 1.
    Dynamic Target Exploration. Temporarily relax decoding to explore a broader output region of the target model and directly sample native harmful responses; select the most harmful as the dynamic target $r^{*}$.
  2. 2.
    Target-Conditioned Optimization. Update $S$ with a small number of gradient optimization iterations to increase the likelihood of producing $r^{*}$ under standard decoding while preserving suffix fluency and avoiding refusal triggers.
Each cycle updates $r^{*}$ through re-sampling, which (i) anchors optimization in high-density regions of the current conditional distribution and (ii) solves the problem of distribution shift induced by optimizing $S$. Formally, the optimization in each cycle solves
$\min_{S}\ \mathcal{L}_{\text{DTA}}(P,S;r^{*})$ $=\mathcal{L}_{\text{resp}}(P,S;r^{*})+\lambda\cdot\mathcal{L}_{\text{suffix}}(S),$ (4)
where $\mathcal{L}_{\text{resp}}$ steers the model toward $r^{*}$ and $\mathcal{L}_{\text{suffix}}$ is a regularizer that promotes fluency while penalizing proximity to refusal-related tokens. The scalar $\lambda>0$ balances these terms.
Figure 2: Overview of DTA. DTA progressively executes an “sampling-optimization cycle” to directly sample the inherent harmful response from the target LLM’s relatively high-probability generation regions and optimize the adversarial suffix. Algorithm 1 shows the details of our DTA.
3.2 Exploration — Dynamically Sampling a Harmful Target
The cornerstone of DTA is that targets are directly sampled from the model’s conditional output distribution, rather than a fixed, low-probability response. Given a harmful prompt $P$ and the current suffix $S$, we query the target model $f$ with a relaxed, high-entropy decoding strategy that increases sampling diversity, thereby directly sampling candidate responses from the target model. Conditioned on P+SP{+}S, we draw $N$ candidates $\{{r_{i}}\}_{i=1}^{N}$ from the target model $f$:
$r_{i}\sim f\big(\cdot\mid P{+}S;\ \text{relaxed decoding}\big),\quad i=1,\ldots,N.$ (5)
Each $r_{i}$ is scored by a harmfulness judge $J$ (e.g., GPTFuzzer), yielding $J(r_{i})\in[0,1]$. We select the most harmful candidate as the cycle’s dynamic target:
$r^{*}=\arg\max_{r_{i}}J(r_{i}).$ (6)
By anchoring optimization to $r^{*}$—a high-density, model-native harmful response—DTA aligns its objective with what the target model already deems plausible, improving its efficiency and stability.
@@ 3.2 Exploration — Dynamically Sampling a Harmful Target
0:  Target LLM $f$, Reference LLM $f_{ref}$, Harmful prompt $P$, Exploration cycles $M$, Optimization iterations per cycle $T$, Samples per cycle NN, Learning rate $\eta$, Harmfulness threshold $\theta_{stop}$
1:  Initialize adversarial suffix $S$ randomly
@@
3:   /* — Phase 1: Sampling — */
@@
4:   Sample NN candidate responses $\{r_{i}\}_{i=1}^{N}$ from $f_{ref}(P+S)$ using high temperature $\tau_{\text{search}}$.
5:   Evaluate harmfulness scores $\{J(r_{i})\}_{i=1}^{N}$ using a judge model.
6:   Select dynamic target response $r^{*}=\arg\max_{r_{i}}J(r_{i})$. {See Eq. 6}
7:   Truncate target to a fixed length for optimization: $r^{*}_{L}$.
8:   /* — Phase 2: Optimization — */
9:   for $t=1$ to $T$ do
10:    // Compute losses based on the target LLM $f$ with standard temperature $\tau_{\text{eval}}$
@@
11:    Compute response loss $\mathcal{L}_{\text{resp}}(P,S;r^{*}_{L})$ over the truncated target. {See Eq. 7}
12:    Compute suffix regularization $\mathcal{L}_{\text{suffix}}(S)=\mathcal{L}_{\text{flu}}(S)-\mathcal{L}_{\text{rej}}(S)$. {See Eqs. 8, 9}
@@
13:    Compute total DTA loss $\mathcal{L}_{\text{DTA}}=\mathcal{L}_{\text{resp}}+\lambda\mathcal{L}_{\text{suffix}}$. {See Eq. 11}
14:    Update suffix via gradient descent: $S\leftarrow S-\eta\cdot\nabla_{S}\mathcal{L}_{\text{DTA}}(P,S)$. {See Eq. 12}
15:    // Check for early stopping condition
@@
16:    Generate a test response $r_{test}=f(P+S)$ under standard decoding.
17:    if $J(r_{test})>\theta_{stop}$ then
18:     break {Terminate refinement early if successful}
19:    end if
20:   end for
21:  end for
@@
22:  Return Optimized adversarial suffix $S^{*}$
3.3 Optimization — Efficiently Steering Toward the Dynamic Target
Given $r^{*}$, DTA updates $S$ via a small number of gradient optimization steps to minimize $\mathcal{L}_{\text{DTA}}$. This objective is composed of two primary components: a response likelihood loss $\mathcal{L}_{\text{resp}}$ and a suffix regularization term $\mathcal{L}_{\text{suffix}}$. Instead of matching the entire, potentially noisy sequence, we truncate $r^{*}$ to a fixed length of response, i.e., $r^{*}_{L}$. We employ this truncation strategy because (1) it mitigates the influence of noisy or incoherent tokens that often appear in the latter part of long generated responses, ensuring that the optimization focuses on more reliable and meaningful content, and (2) it reinforces early-stage control, encouraging the suffix to steer the model output from the very beginning of generation toward harmful behaviors.
The response-likelihood term is then the negative log-likelihood of $r^{*}_{L}$ under standard decoding (temperature $\tau_{\text{eval}}{=}0.7$):
$\mathcal{L}_{\text{resp}}(P,S;r^{*}_{L})=-\sum_{t=1}^{L}\log p_{\tau_{\text{eval}}}(y_{t}\mid y_{<t},,P{+}S).$ (7)
To preserve stealthiness and avoid triggering refusals, we add a suffix regularizer composed of the fluency loss $\mathcal{L}_{flu}$ and the refusal loss $\mathcal{L}_{rej}$:
$\mathcal{L}_{\text{flu}}(S)$ $=-\sum_{j=1}^{|S|}\log p_{\tau_{\text{eval}}}(s_{j}\mid s_{<j}),$ (8)
$\mathcal{L}_{\text{rej}}(S)$ $=\sum_{j=1}^{|S|}\ \sum_{v\in\mathcal{V}_{\text{rej}}}\log p(s_{j}{=}v\mid s{<j}),$ (9)
where $\mathcal{V}_{\text{rej}}$ is a set of common refusal tokens/phrases. We combine them as
$\mathcal{L}_{\text{suffix}}(S)=\mathcal{L}_{\text{flu}}(S)-\mathcal{L}_{\text{rej}}(S),$ (10)

### v3

@@ Abstract
[…]
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide spectrum of natural language tasks (Achiam et al., 2024; Dubey et al., 2024; Yang et al., 2024; Zhao et al., 2024). To steer these capabilities toward beneficial and safe outputs, alignment techniques such as Reinforcement Learning from Human Feedback (RLHF) are employed to instill models with human values (Kirk et al., 2024), which will significantly lower the probability of generating affirmative responses on harmful queries.
Figure 1: DTA targets higher-density harmful responses $T_{\text{sampled}}$ directly sampled from the LLM.
@@
Despite the effectiveness of these alignment techniques, recent work reveals the vulnerability of safety-aligned LLMs to jailbreak attacks (Huang et al., 2025; Zou et al., 2023; Qi et al., 2025; Zeng et al., 2024; Guo et al., 2024; Zhu et al., 2024), where the adversaries aim to elicit harmful content from LLMs by different prompt optimization methods. Among them, most existing gradient-based approaches follow a common paradigm, i.e., optimizing an adversarial suffix $S$ appended to a harmful prompt $P$ to maximize the probability of the target LLM generating a fixed affirmative response $r_{\text{fixed}}$. However, due to the recent advancement in safety alignment, a fixed affirmative response typically lies in an extremely low-density region of the safety-aligned LLM’s conditional output distribution given $P+S$. For instance, in Figure 1, the sequence-level log-likelihood [^1] We report the sum of token-wise log-probabilities over the entire response; long harmful responses naturally yield large-magnitude negative values (e.g., 30-30 to 50-50), even when they are relatively likely under the model. of the fixed affirmative target $r_{\text{fixed}}$ conditional on the initial adversarial prompt is very low (e.g., $\log p(r_{\text{fixed}}\mid P+S)=-52.63$), and even after thousands of optimization iterations, this probability remains low (e.g., $\log p(r_{\text{fixed}}\mid P+S^{*})=-39.07$). This low probability indicates a substantial mismatch between the optimization objective ($r_{\text{fixed}}$) and the target LLM’s output distribution, posing a significant challenge to prompt optimization.
Figure 2: Examples of responses generated by DTA, COLD-Attack, and GCG, and their corresponding target responses. Green text denotes safe responses, and red text denotes harmful responses. GCG and COLD-Attack optimize for thousands of steps toward a fixed harmful response (e.g., “Sure, here is …”), which may induce the target LLM to generate a similar “Sure”-style prefix. However, the full response can still be benign, indicating a failed jailbreak.
Our approach. We propose Dynamic Target Attack (DTA), a new jailbreaking framework that adaptively searches responses over relatively high-density regions of the target LLM’s output space conditioned on the adversarial prompt, and selects the most harmful (yet useful) response as the optimization target, thereby significantly reducing the mismatch between the target and LLM’s output distribution. As shown in Figure 3, DTA (i) uses relaxed decoding with multinomial sampling to directly sample candidate responses from high-density regions of the target LLM $f_{\theta}$’s conditional distribution, and employs a harmfulness judge to select the most harmful one as a temporary target $r^{*}$, (ii) employs a local LLM $f_{\phi}$ to generate a local response $r_{\text{local}}$ under the standard decoding and performs a small number of optimization steps to minimize the mismatch of $r_{\text{local}}$ and $r^{*}$, and (iii) re-samples candidates under the updated adversarial prompt and repeats (i) and (ii) until a stopping criterion is satisfied.
Since the temporary targets reside within the LLM’s relatively high-probability regions, optimizing toward them aligns with the LLM output distribution, thereby easing the optimization process. For example, Figure 1 demonstrates that the probability of the sampled target ($\log p(r_{\text{sampled}}\mid P{+}S)=-36.68$) is much larger than the probability of a fixed target ($\log p(r_{\text{fixed}}\mid P+S)=-52.63$, $\Delta\log p=15.95$). Notably, after a few optimization steps using our DTA, $\log p(r_{\text{sampled}}\mid P{+}S)$ further improves to 28.03-28.03. Since DTA always samples from the updated conditional distribution, which further increases the chance of selecting a more probable target for the next round and thereby further reduces the discrepancy between the target and the LLM’s output distribution. Figure 2 shows that, compared with COLD-Attack and GCG, DTA can sample more harmful and useful target responses, and its induced outputs are correspondingly more harmful.
@@
We demonstrate that DTA effectively jailbreaks multiple recent safety-aligned LLMs under both the white-box and black-box settings. Specifically, under the white-box setting, DTA only requires 200 optimization steps to achieve an average Attack Success Rate (ASR) of over 87% on five recent safety-aligned LLMs, which is over 15% higher than its counterparts. Under the black-box setting, DTA achieves an average ASR of 77.5% with a surrogate model for optimization, which is higher than its counterparts by over 12%. Our primary contributions can be summarized as follows:
@@
  • We introduce Dynamic Target Attack (DTA), a novel jailbreak framework that effectively jailbreaks LLMs under both white-box and black-box settings. In contrast to existing methods, DTA adaptively uses responses over relatively high-density regions as targets, thereby significantly mitigating the mismatch between the target and LLM output distribution.
[…]
  • We further adapt DTA to the black-box setting by optimizing the suffix on a local surrogate model and only using the target API model for sampling the potential target responses. Due to its advantage of sampling the targets from the target API model, DTA can achieve a much higher ASR than other transfer-based attacks.
@@ 2 Background
Figure 3: Overview of DTA. DTA progressively executes an “Exploration-Optimization cycle” to directly sample the inherent harmful response from the target LLM’s relatively high-probability generation regions and optimize the adversarial suffix.
2.1 Paradigms in Gradient-based Jailbreaking
@@
Existing methods, such as GCG (Zou et al., 2023) and COLD-Attack (Guo et al., 2024) are prominent implementations of this paradigm, employing token-level and global-level optimization processes, respectively, to find an effective $S^{*}$.
However, recent researchers have recognized that constraining the optimization to a single, fixed target prefix is overly restrictive and can lead to inefficient optimization. To mitigate this limitation, methods like AdvPrefix (Zhu et al., 2024) are proposed. The core idea of AdvPrefix is to relax the objective by introducing a set of multiple, curated target prefixes, $\mathcal{T}=\{T_{1},T_{2},\cdots,T_{m}\}$. The optimization goal then becomes maximizing the probability of generating any of these targets, often approximated by minimizing the loss of the most likely target in the set:
@@
This multi-target approach provides more flexibility but still relies on a predefined set of external targets. Furthermore, researchers have proposed jailbreaking methods based on Reinforcement Learning (e.g., RLbreaker (Chen et al., 2024)) and adaptive jailbreaking methods (e.g., llm-adaptive-attacks (Andriushchenko et al., 2025). However, all existing methods rely on externally-defined objectives, which may not adaptively align with different target LLMs’ natural harmful responses.
@@ 2.2 Language Model Decoding and Sampling Strategies
In practice, most LLMs adopt greedy sampling or random sampling with a low temperature, which forms a narrow output space. A safety-aligned LLM usually generates refusal responses from this narrow space to harmful prompts. However, recent work indicates that a broader output space randomly sampled from a safety-aligned LLM usually contains harmful responses. Inspired by this observation, we propose DTA to exploit the potentially harmful responses sampled from the LLM as targets to optimize the adversarial suffix, aiming to induce similar responses under the modes of greedy/random sampling with a low temperature.
@@ 3 Methodology
3.1 Problem Formulation
We consider an attacker who jailbreaks a safety-aligned LLM by appending an adversarial suffix $S$ to a harmful prompt $P$. Let $f_{\theta}$ denote the target LLM with vocabulary $\mathcal{V}$ and next-token distribution $p_{\theta}(\cdot\mid x,y_{<t})$. The attacker optimizes a fixed-length suffix $S=(s_{1},\ldots,s_{L_{S}})$ to form the input P+SP{+}S, while keeping model parameters $\theta$ frozen. Given P+SP{+}S, the LLM generates a response $T$ under standard decoding (e.g., temperature $\tau_{\text{eval}}$). We use an external harmfulness judge $J$ that maps $T$ to a score or safe/unsafe label, and deem an attack successful if $T$ is judged unsafe. We consider both white-box and black-box attack settings. In the white-box setting, the attacker has gradient access to the target model $f_{\theta}$ and directly optimizes the suffix $S$ using these gradients. In the black-box setting, the attacker can only query $f_{\theta}$ through an API (without gradients) and thus optimizes $S$ on a local surrogate model $f_{\phi}$, transferring the learned suffix to the target; queries to $f_{\theta}$ are restricted to API-supported decoding controls (e.g., temperature, top-$p$, max tokens). In both settings, for each harmful prompt $P\in\mathcal{D}$, the attacker aims to find a suffix $S$ that maximizes attack success rate (ASR) under a fixed optimization budget (e.g., exploration cycles and optimization steps). Algorithm 1 shows the details of our DTA. All notations are shown in Appendix A.
3.2 Sampling Responses with Relaxed Decoding
The goal of DTA’s exploration stage is to discover harmful candidates that lie in the target LLM’s high-probability output regions under the current adversarial prompt. Rather than optimizing toward a fixed, potentially low-likelihood template, DTA samples candidate responses directly from the model’s conditional distribution conditioned on P+SP{+}S, and then uses a judge to select the most harmful one as a temporary target to optimize the adversarial suffix.
Concretely, we perform stochastic decoding with a non-zero decoding temperature $\tau_{\mathrm{search}}>0$ and multinomial sampling. In practice, we invoke the model with num_beams=1 and do_sample=True, and draw $N$ (approximately) i.i.d. samples per cycle:
$r_{i}\sim f_{\theta}\big(\cdot\mid P{+}S;\,\tau_{\mathrm{search}}\big),\quad i=1,\ldots,N.$ (4)
As each $r_{i}$ is sampled from $p_{\theta}(\cdot\mid P{+}S)$, the candidate set concentrates on responses with non-negligible probability mass under the current prompt. We use a judge $J$ to evaluate harmfulness scores $J(r_{i})\in[0,1]$ and select
$r^{*}=\arg\max_{r_{i}}J(r_{i})$ (5)
as a temporary target that is both harmful and useful, which substantially eases the subsequent optimization process. Then, we truncate $r^{*}$ to a fixed-length prefix $r^{*}_{L}$ before optimization. This truncation could mitigate noise in the tail of long generations and reinforces early-stage control: the suffix is trained to steer the model from the beginning of the response toward harmful behavior.
In the black-box setting, DTA only requires query access to the target API and samples candidates directly from the target LLM’s conditional distribution under P+SP{+}S. Therefore, the selected dynamic target $r^{*}$ is more natural and higher-probability than the fixed target response, which often lies in low-likelihood regions of the target model.
Across exploration–optimization cycles, updating $S$ gradually shifts the conditional distribution away from refusal modes and increases the probability mass assigned to harmful regions. As a result, even if early cycles only surface benign or borderline candidates, repeated multinomial sampling under the updated prompt makes harmful candidates increasingly likely to appear in later cycles. We provide a brief discussion of this behavior in Appendix B.
@@ 3.2 Sampling Responses with Relaxed Decoding
0:  Target LLM $f_{\theta}$, harmful prompt $P$, judge model $J$, exploration cycles $M$, optimization steps per cycle $T$, samples per cycle NN, learning rate $\eta$, exploration temperature $\tau_{\text{search}}$, evaluation temperature $\tau_{\text{eval}}$, harmfulness threshold $\Theta_{\text{stop}}$
1:  Initialize adversarial suffix $S$ (and its continuous parameters, e.g., logits $\mathbf{Z}$) randomly
@@
3:  /* — Phase 1: Dynamic target exploration — */
@@
4:  Sample NN candidate responses $\{r_{i}\}_{i=1}^{N}$ from $f_{\theta}(P{+}S)$ using relaxed decoding with temperature $\tau_{\text{search}}$ (multinomial sampling) {See Sec. 3.2}
5:  Select dynamic target response $r^{*}$ {See. Eq. 5}
6:  Truncate $r^{*}$ to length $L$ to obtain the target prefix $r^{*}_{L}$
7:  /* — Phase 2: Target-conditioned optimization — */
8:  for $t=1$ to $T$ do
@@
9:   Compute response loss $\mathcal{L}_{\text{resp}}(P,S;r^{*}_{L})$ under standard decoding with $\tau_{\text{eval}}$ {See Eq. 8}
@@
10:   Form total loss $\mathcal{L}_{\text{DTA}}=\mathcal{L}_{\text{resp}}+\lambda\,\mathcal{L}_{\text{suffix}}$ {See Eq. 12}
11:   Project $\mathbf{Z}$ to a discrete suffix $S$ {See Sec. 3.3}
12:   // Early-stopping check under evaluation decoding
@@
13:   Generate a test response $r_{\text{test}}$ from $f_{\theta}(P{+}S)$ using standard decoding with $\tau_{\text{eval}}$
14:   if $J(r_{\text{test}})>\Theta_{\text{stop}}$ then
@@
15:     return optimized adversarial suffix $S^{*}\leftarrow S$ {Terminate refinement early if successful}
16:   end if
17:  end for
18:  end for
19:  return final adversarial suffix $S^{*}\leftarrow S$
3.3 Optimization Procedure
Our optimization procedure follows the continuous logit-parameterization paradigm introduced by COLD-Attack (Guo et al., 2024): the parameters of the target LLM $f_{\theta}$ are frozen, and we only optimize a continuous representation of the suffix while periodically projecting it back to discrete tokens. Notably, in the black-box setting, we employ a local LLM $f_{\phi}$ as a surrogate to back-propagate gradients and optimize suffix; In the white-box setting, we directly use the target LLM $f_{\theta}$ (i.e., $f_{\phi}=f_{\theta}$). Below, we describe this procedure in detail.
Continuous parameterization of the suffix.
Let the suffix length be $L_{S}$ and the vocabulary size be $|\mathcal{V}|$. Instead of directly optimizing discrete tokens, DTA maintains a trainable logit matrix $\mathbf{Z}\in\mathbb{R}^{L_{S}\times|\mathcal{V}|}$, where the $i$-th row $\mathbf{z}_{i}$ represents the logits over the vocabulary at position $i$. We obtain a relaxed token distribution at each position by applying a low-temperature softmax:
$\boldsymbol{\pi}_{i}=\mathrm{softmax}\!\left(\frac{\mathbf{z}_{i}}{\tau_{\mathrm{opt}}}\right),$ (6)
where we set $\tau_{\mathrm{opt}}=10^{-3}$ as a very small optimization temperature, so that the softmax distribution becomes nearly one-hot and thus closely approximates discrete token selections while remaining differentiable. Given the frozen token embedding matrix $\mathbf{E}\in\mathbb{R}^{|\mathcal{V}|\times d}$, the embedding for position $i$ in the suffix is then the expectation
$\mathbf{e}_{i}=\boldsymbol{\pi}_{i}^{\top}\mathbf{E}.$ (7)
Stacking $\{\mathbf{e}_{i}\}_{i=1}^{L_{S}}$ yields a continuous embedding sequence for the suffix $S$, which is concatenated with the embedding of the prompt $P$ and fed into the LLM.
Loss function.
For a fixed harmful prompt $P$ and its current target prefix $r^{*}_{L}=(t^{*}_{0},\ldots,t^{*}_{L-1})$, DTA minimizes a differentiable objective that encourages the model to generate $r^{*}_{L}$ when conditioned on P+SP{+}S. Let $f_{\theta}(P{+}S)$ denote the sequence of output logits of the target LLM and $p_{\theta}$ the corresponding token distribution after softmax. We define the response loss as the cross-entropy between the model’s predictive distribution and the target tokens:
$\mathcal{L}_{\mathrm{resp}}(P,S;r^{*}_{L})=-\frac{1}{L}\sum_{j=0}^{L-1}\log p_{\theta}\!\left(t^{*}_{j}\mid P{+}S,t^{*}_{<j}\right).$ (8)
Suffix regularization.
To avoid degenerate suffixes and to encourage fluent, non-refusal text, we instantiate the regularizer $\mathcal{L}_{\mathrm{suffix}}(S)$ as a combination of a fluency term and a refusal-penalty term. Let $S=(s_{1},\ldots,s_{L_{S}})$ and let $p_{\theta}$ denote the token distribution of the target LLM $f_{\theta}$. We define the fluency loss as the negative log-likelihood of the suffix under the model:
$\mathcal{L}_{\mathrm{flu}}(S)=-\sum_{j=1}^{L_{S}}\log p_{\theta}\big(s_{j}\mid s_{<j}\big),$ (9)
which encourages $S$ to lie in a high-probability, on-distribution region. We also construct a small set $\mathcal{V}_{\mathrm{rej}}$ of common refusal tokens or phrases (e.g., “cannot help with”, “as an AI model”) and penalize placing probability mass on them:
$\mathcal{L}_{\mathrm{rej}}(S)=\sum_{j=1}^{L_{S}}\sum_{v\in\mathcal{V}_{\mathrm{rej}}}p_{\theta}\big(s_{j}{=}v\mid s_{<j}\big).$ (10)
The overall regularizer is then
$\mathcal{L}_{\mathrm{suffix}}(S)=\mathcal{L}_{\mathrm{flu}}(S)-\mathcal{L}_{\mathrm{rej}}(S),$ (11)